# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 8.2 Posterior Performance Analysis of SNNs
In Figure 6, we compare the mean and covariance of various SNNs (obtained via Monte Carlo sampling using $41e$ 4 samples) and the relative GMM approximation on a set of $30$ randomly collected input points for various neural network architectures on the MNIST and CIFAR-10 datasets. We observe that the GMMs match the MC approximation of the true mean and kernel. However, unlike the GMM approximations, the MC approximations lack formal guarantees of correctness, and additionally, computing the MC approximations is generally two orders of magnitudes slower than computing the GMM approximations. By analyzing Figure 6, it is possible to observe that since the architectures allow for the training of highly accurate networks, each row in the posterior mean reflects that the classes are correctly classified. For the networks trained using VI, the kernel matrix clearly shows the correlations trained by the SNN. For the dropout network, the GMM visualizes the limitations of dropout learning in capturing correlation among data points. This finding aligns with the result in Gal and Ghahramani (2016), which shows that Bayesian inference via dropout is equivalent to VI inference with an isotropic Gaussian over the weights per layer.
@@
Figure 7: Visualization of the prior distribution of (a) a SNN with an isotropic Gaussian distribution over its weights; (b) a zero mean GP with the RBF kernel; (c-d) SNNs with a Gaussian distribution over its weight that is optimized to mimic the GP in (b) via, respectively, our framework and the approach proposed in Tran et al. (2022). The SNN has 2 hidden layers, 128 neurons per layer, and the tanh activation function, while the GP has zero mean and a RBF kernel with length-scale $0.5$ and signal variance $1$. The red lines, blue shaded region, and blue lines respectively indicate the mean, $99.7\%$ credible intervals, and a subset of the MC samples from the output distributions.
@@ 8.3 Functional priors for neural networks via GMM approximation
In this section, to highlight another application of our framework, we show how our results can be used to encode functional priors in SNNs. We start from Figure 7(a), where we plot functions sampled from a two hidden layer SNN with an isotropic Gaussian distribution over its weights, which is generally the default choice for priors in Bayesian learning (Fortuin, 2022). As it is possible to observe in the figure, the sampled functions tend to form straight horizontal lines; this is a well-known pathology of neural networks, which becomes more and more exacerbated with deep architectures (Duvenaud et al., 2014), Instead, one may want to consider a prior model that reflects certain characteristics that we expect from the underlying true function, e.g., smoothness or oscillatory behavior. Such characteristics are commonly encoded via a Gaussian process with a given mean and kernel (Tran et al., 2022), as shown for instance in Figure 7(b), where we plot samples from a GP with zero mean and the RBF kernel (MacKay, 1996). We now demonstrate how our framework can be used to encode an informative functional prior, represented as a GP, into a SNN. In particular, in what follows, we assume centered Mean-Field Gaussian priors on the weights, i.e., $p_{{w}}=\mathcal{N}({{\bar{0}}},\mathrm{diag}(\psi))$, where $\psi$ is a vector of possibly different parameters representing the variance of each weight. Our objective is to optimize $\psi$ to minimize the 2-Wasserstein distance between the distributions of the GP and the SNN at a finite set of input points, which we do by following the approach described in Section 7.2. For the low-dimensional 1D regression problem, the input points contain uniform samples in the input domain. For the high-dimensional UCI problems, we use the training sets augmented with noisy samples. We solve the optimization problem as described in Section 7.2, setting $\beta$ to $0.01$, using mini-batch gradient descent with randomly sampled batches.
@@
Table 4: The empirical estimates of the relative 2-Wasserstein distance between the prior distributions of zero-mean GPs with the RBF kernel and SNNs with uninformative or GP-induced priors over the weights, at a subset of $20$ points from the test set. The GP-induced prior are either obtained via Algorithm 1 with signatures of size $10$ and compression size (M) of $1$ as described in Section 7.2, or via the method in Tran et al. (2022).
@@
We continue our analysis on several UCI regression datasets to investigate whether SNNs with an informative prior lead to improved posterior performance compared to uninformative priors. We evaluate the impact of an informative prior on the predictive performance of SNNs using the root-mean-square error (RMSE) and negative log-likelihood (NLL) metrics. The RMSE solely measures the accuracy of the mean predictions, whereas the NLL evaluates whether the predicted distribution matches the actual distribution of the outcomes, taking into account the uncertainty of the predictions. Lower RMSE and NLL values indicate better performance. Figure 8 shows that for all datasets, the GP outperforms the SNN with an uninformative prior both in terms of mean and uncertainty accuracy. By encoding the GP prior in the SNN, we can greatly improve the predictive performance of the SNN, bringing it closer to the performance of the GP. The remaining performance gap to the GP can be explained by the VI error in approximating the posterior distribution.
Figure 8: Report of the predictive performance of a GP with zero-mean and RBF kernel, and a $[2\times 128]$ SNN with tanh activation function on several UCI regression datasets in terms of the root-mean-squared error (RMSE) and negative log-likelihood (NLL). The dots and bars represent the means and standard deviations over the 10 random test splits of the datasets, respectively. The results for the GP are shown in blue, the results for the SNN with an uninformative isotropic Gaussian prior on the weights are in yellow, and the results for the SNN with the GP-induced informative priors obtained via our approach are in green.
@@ 9 Conclusion
We introduced a novel algorithmic framework to approximate the input-output distribution of a SNN with a GMM, providing bounds on the approximation error in terms of the Wasserstein distance. Our framework relies on techniques from quantization theory and optimal transport, and is based on a novel approximation scheme for GMMs using signature approximations. We performed a detailed theoretical analysis of the error introduced by our GMM approximations and complemented the theoretical analysis with extensive empirical validation, showing the efficacy of our methods in both regression and classification tasks, and for various applications, including prior selection and uncertainty quantification. We should stress that in this paper we focused on finding GMM approximations of neural networks at a finite set of input points, thus also accounting for the correlations among these points. While reasoning for a finite set of input points is standard in stochastic approximations (Yaida, 2020; Basteri and Trevisan, 2024) and has many applications, as also highlighted in the paper, for some other applications, such as adversarial robustness of Bayesian neural networks (Wicker et al., 2020), one may require approximations that are valid for compact sets of input points. We believe this is an interesting future research question.
Appendix A Proofs
@@
We present the proofs for all results discussed in the main text of this work. These proofs rely on several less commonly reported properties of the Wasserstein distance, which we will first discuss in Section A.1. After this, we present the proofs ordered per section.
A.1 Properties of the Wasserstein Distance
Lemma 23 (2-Wasserstein Distance for Independent Joint Distributions)

### v2

@@ 8.2 Posterior Performance Analysis of SNNs
In Figure 7, we compare the mean and covariance of various SNNs (obtained via Monte Carlo sampling using $1e4$ samples) and the relative GMM approximation on a set of $30$ randomly collected input points for various neural network architectures on the MNIST and CIFAR-10 datasets. We observe that the GMMs match the MC approximation of the true mean and kernel. However, unlike the GMM approximations, the MC approximations lack formal guarantees of correctness, and additionally, computing the MC approximations is generally two orders of magnitudes slower than computing the GMM approximations. By analyzing Figure 7, it is possible to observe that since the architectures allow for the training of highly accurate networks, each row in the posterior mean reflects that the classes are correctly classified. For the networks trained using VI, the kernel matrix clearly shows the correlations trained by the SNN. For the dropout network, the GMM visualizes the limitations of dropout learning in capturing correlation among data points. This finding aligns with the result in Gal and Ghahramani (2016), which shows that Bayesian inference via dropout is equivalent to VI inference with an isotropic Gaussian over the parameters per layer.
@@
Figure 8: Visualization of the prior distribution of (a) a SNN with an isotropic Gaussian distribution over its parameters; (b) a zero mean GP with the RBF kernel; (c-d) SNNs with a Gaussian distribution over its weight that is optimized to mimic the GP in (b) via, respectively, our framework and the approach proposed in Tran et al. (2022). The SNN has 2 hidden layers, 128 neurons per layer, and the tanh activation function, while the GP has zero mean and a RBF kernel with length-scale $0.5$ and signal variance 11. Red lines, blue shading, and blue lines respectively indicate the mean, $99.7\%$ credible intervals, and a subset of the MC samples from the output distributions.
@@ 8.3 Functional priors for neural networks via GMM approximation
In this section, to highlight another application of our framework, we show how our results can be used to encode functional priors in SNNs. We start from Figure 8(a), where we plot functions sampled from a two hidden layer SNN with an isotropic Gaussian distribution over its parameters, which is generally the default choice for priors in Bayesian learning (Fortuin, 2022). As it is possible to observe in the figure, the sampled functions tend to form straight horizontal lines; this is a well-known pathology of neural networks, which becomes more and more exacerbated with deep architectures (Duvenaud et al., 2014), Instead, one may want to consider a prior model that reflects certain characteristics that we expect from the underlying true function, e.g., smoothness or oscillatory behavior. Such characteristics are commonly encoded via a Gaussian process with a given mean and kernel (Tran et al., 2022), as shown for instance in Figure 8(b), where we plot samples from a GP with zero mean and the RBF kernel (MacKay, 1996). We now demonstrate how our framework can be used to encode an informative functional prior, represented as a GP, into a SNN. In particular, in what follows, we assume centered Mean-Field Gaussian priors on the parameters, i.e., $p_{{w}}=\mathcal{N}({{\bar{0}}},\mathrm{diag}(\psi))$, where $\psi$ is a vector of possibly different parameters representing the variance of each weight. Our objective is to optimize $\psi$ to minimize the 2-Wasserstein distance between the distributions of the GP and the SNN at a finite set of input points, which we do by following the approach described in Section 7.2. For the low-dimensional 1D regression problem, we use uniformly sampled input points. For the high-dimensional UCI problems, we use the training sets augmented with noisy samples. We solve the optimization problem as described in Section 7.2, setting $\beta$ to $0.01$, using mini-batch gradient descent with randomly sampled batches.
@@
Table 4: The empirical estimates of the relative 2-Wasserstein distance between the prior distributions of zero-mean GPs with the RBF kernel and SNNs with uninformative or GP-induced priors over the parameters, at a subset of $20$ points from the test set. The GP-induced prior are either obtained via Algorithm 1 with signatures of size $10$ and compression size ($M_{k}$) of $1$ for all layers as described in Section 7.2, or via the method in Tran et al. (2022).
Figure 9: Report of the predictive performance of a GP with zero-mean and RBF kernel, and a $2\times 128]$ SNN with tanh activation function on several UCI regression datasets in terms of the root-mean-squared error (RMSE) and negative log-likelihood (NLL). The dots and bars represent the means and standard deviations over the 10 random test splits of the datasets, respectively. The results for the GP are shown in blue, the results for the SNN with an uninformative isotropic Gaussian prior on the parameters are in yellow, and the results for the SNN with the GP-induced informative priors obtained via our approach are in green.
@@
We continue our analysis on several UCI regression datasets to investigate whether SNNs with an informative prior lead to improved posterior performance compared to uninformative priors. We evaluate the impact of an informative prior on the predictive performance of SNNs using the root-mean-square error (RMSE) and negative log-likelihood (NLL) metrics. The RMSE solely measures the accuracy of the mean predictions, whereas the NLL evaluates whether the predicted distribution matches the actual distribution of the outcomes, taking into account the uncertainty of the predictions. Lower RMSE and NLL values indicate better performance. Figure 9 shows that for all datasets, the GP outperforms the SNN with an uninformative prior both in terms of mean and uncertainty accuracy. Encoding the GP prior in the SNN greatly improves its predictive performance, bringing it closer to that of the GP, with the remaining gap explained by the VI approximation error of the posterior distribution.
@@ 9 Conclusion
We introduced a novel algorithmic framework to approximate the input-output distribution of a SNN with a GMM, providing bounds on the approximation error in terms of the Wasserstein distance. Our framework relies on techniques from quantization theory and optimal transport, and is based on a novel approximation scheme for GMMs using signature approximations. We performed a detailed theoretical analysis of the error introduced by our GMM approximations and complemented the theoretical analysis with extensive empirical validation, showing the efficacy of our methods in both regression and classification tasks, and for various applications, including prior selection and uncertainty quantification. An interesting future direction is to employ the framework in safety-critical control, planning, or decision-making applications, where our method critically enables principled uncertainty quantification. We should stress that in this paper we focused on finding GMM approximations of neural networks at a finite set of input points, thus also accounting for the correlations among these points. While reasoning for a finite set of input points is standard in stochastic approximations (Yaida, 2020; Basteri and Trevisan, 2024) and has many applications, for some other applications, such as adversarial robustness of Bayesian neural networks (Wicker et al., 2020), one may require approximations that are valid for compact sets of input points. We believe this is an interesting future research question.
Appendix A Properties of the Wasserstein Distance
@@
The results in this work rely on several less commonly reported properties of the Wasserstein distance, which are recalled below.
Lemma 28 (Closeness in ${\mathbb{W}}_{\rho}$ Implies Closeness in $\rho$-Moments)
For $p,q\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$ and $\rho\in\mathbb{N}$, it holds that
$\left|{\mathbb{E}}_{{{x}}\sim p}[\|{{x}}\|^{\rho}]^{1/\rho}-{\mathbb{E}}_{{{z}}\sim q}[\|{{z}}\|^{\rho}]^{1/\rho}\right|\leq{\mathbb{W}}_{\rho}(p,q).$ (36)
Proof  The result follows directly from Proposition 7.29 in Villani et al. (2009), applied to the $1$-Lipschitz function $\phi({{x}})=\|{{x}}\|$.
Lemma 29 (Translation and Rotation Invariance of ${\mathbb{W}}_{\rho}$)
Let $\tau:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ be an isometry of ${\mathbb{R}}^{n}$, i.e., $\tau({{x}})={{R}}{{x}}+{{b}}$ for some orthogonal matrix ${{R}}\in{\mathbb{R}}^{n\times n}$ and vector ${{b}}\in{\mathbb{R}}^{n}$. Then, for $\rho\in\mathbb{N}$ and $p,q\in{\mathcal{P}}({\mathbb{R}}^{n})$, we have
${\mathbb{W}}_{\rho}(\tau\#p,\tau\#q)={\mathbb{W}}_{\rho}(p,q).$ (37)
Proof  Since $\tau$ preserves the Euclidean norm, $\|\tau({{x}})-\tau({{z}})\|=\|{{x}}-{{z}}\|$, any coupling $\gamma\in\Gamma(p,q)$ is mapped to a coupling $(\tau\times\tau)\#\gamma\in\Gamma(\tau\#p,\tau\#q)$ with the same transport cost. The optimal value is therefore unchanged.
Lemma 30 (${\mathbb{W}}_{2}$ for Product Distributions)